OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the second equation is substituted into the first equation. 2x – 7y = 4 3x + y = –17

OpenStudy (anonymous):

A.) 2(3x – 17) – 7y = 4 B.) 2(–3x – 17) – 7y = 4 C.) 2x – 7(3x – 17) = 4 D.) 2x – 7(–3x – 17) = 4

OpenStudy (anonymous):

I have gotten the answer 2x-7(3x-17)=4

OpenStudy (anonymous):

not sure if im correct though